im not quit sure what cat. this question would go under so i tried to find the closes one...i apologize if it is off topic...
a 40 ft. board is to be cut into two peices. The LONGER PIECE is to be 7 feet longer than twice the SHORTER peice. How long is each peice?
let x= shorter peice
___=the longer peice
equation:

Answer by LilSkittleMd(119)

x=shorter piece
y=longer piece
The 40 ft board is to be cut into 2 pieces,which is the equation:
x+y=40
The Longer piece is 7 feet longer than twice the shorter piece, its equation is:
y=7+2x
Then you substitute the y= equation into the x+y equation:
x+7+2x=40 (add together the x's, subtract 7 from 40)
3x=33 (divide by 3 to isolate x)
x=11
Then plug x into either equation
y=7+2(11)
y=7+22
y=29
The shorter piece is 11 feet and the longer piece is 29 feet.
Check:
11+29=40
40=40
29=7+2(11)
29=7+22
29=29